# Fully Open Source Moxin-LLM Technical Report

### v2

@@ 4.1 Evaluation Tasks
4.2 Evaluation Results
We name the initial model as Moxin-7B-original, which presents the foundation model before fine-tuning on the training data of the evaluation datasets. After subsequent partial fine-tuning of Moxin-7B-original on the training data of the evaluation datasets, we developed Moxin-7B-finetuned, enabling direct assessment of how targeted fine-tuning affects model performance.
@@ 4.2.2 Few-Shot Evaluation
4.3 Alignment Evaluation
Table 2: Performance for various chat models.
Model MTbench
Moxin Chat 7B 6.42
Llama 2 Chat 13B 6.65
Vicuna 13B 6.57
Llama 2 Chat 7B 6.27
Vicuna 7B 6.17
Alpaca 13B 4.53
We evaluate the alignment performance on MTBench [80]. It is a two-round conversation dataset with 80 questions. It covers eight dimensions (reasoning, roleplay, math, coding, writing, humanities, STEM, and information extraction) with 10 questions for each dimension. The model needs to answer the first question and then refine its previous response following additional specific instructions. We use GPT-4 as a judge model to provide scores (between 1-10) for the quality of responses. Our Moxin-7B-chat achieves superior performance on MTbench compared with baselines, as shown in Table 2.
@@ 4.4 Generation Examples
We demonstrate some text generation examples from our Moxin Chat 7B and Llama2 Chat 7B in Table 4.4 and 4.4.
@@ 5 Conclusion
The field of Large Language Models has witnessed a significant shift toward open-source development, fostering innovation within the AI community. However, a critical challenge emerges: many purportedly open-source models withhold essential components necessary for full understanding and reproducibility, creating barriers that limit both academic advancement and commercial adoption. This does not not only hamper scientific progress but also prevent businesses from fully leveraging these models for innovative applications, ultimately diminishing potential societal benefits and economic value creation. To address these limitations, we introduce Moxin 7B, a fully open-source language model developed in accordance with the Model Openness Framework (MOF), providing comprehensive access to pre-training code, configurations, training and fine-tuning datasets, and all intermediate checkpoints. Our evaluation results demonstrate that the Moxin 7B achieves superior zero-shot evaluation results compared to popular 7B models while maintaining competitive few-shot capabilities. We wish to see more work that establishes new standard for reproducible research in language model development, fostering a more inclusive and economically vibrant AI ecosystem.

### v3

@@ 6.1 Evaluation Tasks
6.2 Pre-Training Evaluation
We name the initial model as Moxin-7B-original, which presents the foundation model before fine-tuning on the training data of the evaluation datasets. After subsequent partial fine-tuning of Moxin-7B-original on the training data of the evaluation datasets, we developed Moxin-7B-finetuned, enabling direct assessment of how targeted fine-tuning affects model performance. We release our Moxin-7B-finetuned model as Moxin 7B Base model.
@@ 6.2.2 Few-Shot Evaluation
6.3 Post-Training Evaluation
In our post-training, with Tülu 3, we further fine-tune our base model with SFT and then DPO, leading to our SFT model and DPO model. We also adopt the Infinity Instruct dataset [63] to train another version of our instruct model based on our base model, and release this instruct model as Moxin-7B-II. We demonstrate the performance in Table 6.3 for zero-shot evaluation and Table 6.3 for few show evaluation. We also use the OLMES framework from Tulu 3 [62] to evaluate our model across a range of tasks and show the results in Table 6.3. We can observe that our DPO model can achieve comparable performance with other SOTA instruct models. We release our DPO model as Moxin 7B Instruct model.
[t] Performance comparison for various models in zero-shot evaluation. Models HellaSwag WinoGrade PIQA ARC-E ARC-C Ave Mistral 8B Instruct 79.08 73.56 82.26 79.88 56.57 74.27 Llama3.1 8B Instruct 79.21 74.19 80.79 79.71 55.03 73.79 Qwen2.5 7B Instruct 80.5 71.03 80.47 81.31 55.12 73.69 Moxin - 7B - II 79.32 72.93 81.56 80.43 56.91 74.23 Moxin - 7B - SFT 81.44 73.09 81.07 79.8 54.67 74.01 Moxin - 7B - DPO 85.7 73.24 81.56 81.1 58.02 75.92
[t] Performance comparison for various models in few-shot evaluation. Model ARC-C Hellaswag MMLU WinoGrade Ave Mistral 8B Instruct 62.63 80.61 64.16 79.08 71.62 Llama3.1 8B Instruct 60.32 80 68.18 77.27 71.44 Qwen2.5 7B Instruct 66.72 81.54 71.3 74.59 73.54 Moxin - 7B - II 61.35 82.1 62.95 77.98 71.095 Moxin - 7B - SFT 60.11 83.43 60.56 77.56 70.42 Moxin - 7B - DPO 64.76 87.19 58.36 76.32 71.66
[t] Performance comparison for various models in olmes evaluation. Models/Datasets GSM8K MATH Humaneval Humaneval plus MMLU PopQA BBH TruthfulQA Ave Qwen2.5 7B Instruct 83.8 14.8 93.1 89.7 76.6 18.1 21.7 63.1 57.61 Gemma2 9B Instruct 79.7 29.8 71.7 67 74.6 28.3 2.5 61.4 51.88 Moxin - 7B - II 71.04 21 78.21 72.35 63.27 27.98 44.33 56.22 54.42 Moxin - 7B - DPO 81.19 36.42 82.86 77.18 60.85 23.85 57.44 55.27 59.38
6.4 CoT Evaluation
Based on our instruct model, we further finetune the model with reasoning data based on SFT and then train using reinforcement learning with GRPO following DeepSeek R1. Our final model is trained with DeepScaleR [67] and released as Moxin 7B Reasoning model. We also trained another version of the reasoning model with AReal [68] starting from our instruct model, with its name as Moxin - 7B - Reasoning - AReal. We evaluate our reasoning models on competition-level mathematics benchmarks, including AMC 2023, MATH-500, Minerva Math, and OlympiadBench. We report the Pass@1 accuracy. We use Qwen Math’s codebase for evaluation following [109, 110] and show the comparison in Table 6.4. We can observe that Moxin Reasoning model can achieve outstanding performance compared with baselines, demonstrating the effectiveness of RL for small LLMs such as 7B models. Our Moxin Reasoning model trained with DeepScaleR performs better than that trained with AReal.
[t] Performance comparison for various models on reasoning evaluation. Models/Datasets MATH 500 AMC Minerva Math OlympiadBench Ave Qwen2.5-Math-7B-Base 52.4 52.5 12.9 16.4 33.55 Qwen2.5-Math-7B-Base + 8K MATH SFT 54.6 22.5 32.7 19.6 32.35 Llama-3.1-70B-Instruct 64.6 30.1 35.3 31.9 40.48 Moxin - 7B - Reasoning - AReal 68.6 50 16.9 31.7 41.8 Moxin - 7B - Reasoning 68 57.5 16.9 30.4 43.2
@@ 6.5 Generation Examples
We demonstrate some text generation examples from our Moxin Chat 7B and Llama2 Chat 7B in Table 6.5 and 6.5. Moxin Chat 7B is developed by finetune the Moxin base model on the Tulu2 dataset. The response of Moxin Chat 7B is more accurate and concise.
@@ 7 Conclusion
The field of Large Language Models has witnessed a significant shift toward open-source development, fostering innovation within the AI community. However, a critical challenge emerges: many purportedly open-source models withhold essential components necessary for full understanding and reproducibility, creating barriers that limit both academic advancement and commercial adoption. This not only hampers scientific progress, but also prevents businesses from fully leveraging these models for innovative applications, ultimately diminishing potential societal benefits and economic value creation. To address these limitations, we introduce Moxin 7B, a fully open-source language model developed in accordance with the Model Openness Framework (MOF), providing comprehensive access to pre-training code, configurations, training and fine-tuning datasets, and all intermediate checkpoints. We further finetune the Moxin Base model with open-source post-training frameworks and instruction/CoT data to obtain Moxin Instruct and Moxin Reasoning models. Experiments show that our models achieve superior performance in various evaluations such as zero-shot, few-shot, and CoT evaluations. We wish to see more work that establishes new standards for reproducible research in language model development, fostering a more inclusive and economically vibrant AI ecosystem.